OpenStudy (anonymous):

Which proportion models this problem? On a soccer team, 8 out of 19 players surveyed say they had two or more siblings. The league has 570 players. Which is the best prediction of the number of players in the league that have two or more siblings? A. x/8=19/570 B.x/570=19/8 C.x/570=8/19 D.x/19=8/570

OpenStudy (anonymous):

@thecollinjc1 do u know this to? lol ill give ya a medal :)

OpenStudy (anonymous):

8 of 19 players said that they had 2 or more siblings. 8 out of 19 is equal to what out of 570 8/19 = x/570 This proportion can be written in different ways. 8/19 = x/570 cross multiplies to 19*x = 8*570 The task is to find a proportion among the answer options that cross multiplies to the same equation. Option C is x/570 = 8/19. That cross multiplies to 19*x = 8*570 and is the correct answer.